OpenStudy (anonymous):

If f(x) = x + 4 ,find f(-7).

OpenStudy (anonymous):

plug -7 in for x f(-7) = (-7) + 4 = -3

OpenStudy (anonymous):

f(x)=x + 4 f(-7)= -7+4 f(-7)= -3

OpenStudy (anonymous):

Alright so we have f(-7) and whenever you see a number that is representing a variable by f, it literally means to plug in the number to represent the variable. So, take x out of the picture and plug it in for x. Once you do that, your new equation will be -7+4 and after simplifying that, your answer is then -3